OpenStudy (anonymous):

graph using slope and the y intercept x+2y=8

OpenStudy (anonymous):

Rewrite making y the subject of the equation x+2y=8 2y = 8 -x (subtracting x from both sides) Now divide both sides by 2 y = 4 -x/2 Slope = -1/2 y-intercept 4

OpenStudy (anonymous):